# Online Flow Time Minimization: Tight Bounds for Non-Preemptive Algorithms

### v2

@@ 1.4 Paper Organization
The remainder of this paper is organized as follows. In Section 3, we introduce our core technical contribution: the online rank-based partitioning framework that serves as the foundation for all our algorithms. The analysis of the NSJF algorithm, used for handling small jobs within this framework, is detailed in Appendix A.
We then present our main algorithmic results. Section 4 develops our $O(\sqrt{n/m})$-competitive randomized non-preemptive algorithm. Section 5 presents our deterministic non-preemptive algorithm. Section 6 details our deterministic $O(\sqrt{n/m})$-competitive algorithm using kill-and-restart for $m\geq 2$. We then address the unknown-$n$ setting in Section 7 with an extension of this algorithm.
@@ 1.5 Related Works
In the classic non-preemptive offline setting, the problem is known to be NP-hard. A key result by DBLP:journals/jcss/LeonardiR07 provided an $O(\sqrt{n/m}\log(n/m))$-approximation algorithm, which stood as the best-known guarantee for many years. On the approximation hardness front, DBLP:journals/siamcomp/KellererTW99 established a polynomial-time lower bound of $\Omega(\sqrt{n})$ for the single-machine case, later complemented by DBLP:journals/jcss/LeonardiR07, presenting an $\Omega(n^{1/3-\varepsilon})$ lower bound for multiple machines when $n\geq m^{4/\varepsilon}$. In the online setting, the problem is notoriously difficult, with a simple greedy algorithm achieving a $\Theta(n/m)$ competitive ratio, which for a long time was the best-known result for deterministic algorithms.
@@
To circumvent the hardness of non-preemptive scheduling, a significant body of work has focused on the preemptive model, where jobs can be suspended and resumed. In this setting, the folklore SRPT (Shortest Remaining Processing Time) policy is optimal on a single machine. For multiple machines, SRPT achieves a tight competitive ratio of $\Theta(\log(n/m))$ when job migration is allowed [DBLP:journals/jcss/LeonardiR07]. For the more restrictive model without migration, DBLP:journals/siamcomp/AwerbuchALR02 presented an algorithm with a competitive ratio of $O(\log(\min\{\log n,\log P\}))$.
@@
Another prominent line of research has explored models that grant additional power to the online non-preemptive algorithm. One common approach is resource augmentation, where the algorithm is provided with faster or more machines than the adversary. DBLP:journals/algorithmica/PhillipsSTW02 showed that an $O(1)$-competitive ratio is achievable with $m\log P$ machines, where $P$ is the max-to-min job size ratio. Another popular relaxation is the rejection model, where the algorithm can choose to reject a small fraction of jobs. DBLP:journals/jcss/ChoudhuryDGK18 first introduced this notion for flow time, with subsequent works [DBLP:conf/icalp/GuptaKL18, DBLP:conf/esa/LucarelliMTST18] achieving an $O(1/\varepsilon^{3})$-competitive algorithm for rejecting an $\varepsilon$-fraction of jobs.
@@
The kill-and-restart capability, introduced by DBLP:journals/siamcomp/ShmoysWW95 in the context of makespan minimization, represents a minimal form of preemption. This model has been successfully applied to other objectives, demonstrating its power to surpass non-preemptive lower bounds. For minimizing total completion time, DBLP:journals/jal/SteeP05 designed a $3/2$-competitive algorithm, breaking the non-preemptive barrier of $1.582$. Even more dramatically, for non-clairvoyant scheduling, DBLP:journals/mp/JagerSWW25 showed that kill-and-restart reduces the competitive ratio for total completion time from $\Omega(n)$ to $O(1)$. Despite these successes, the potential of kill-and-restart for flow time minimization has remained largely unexplored. The primary work in this specific area was by DBLP:journals/tcs/EpsteinS03, who established several lower bounds, including an $\Omega(\sqrt{n})$ bound for randomized algorithms on a single machine, but no upper bounds were known.
@@ 2 Preliminaries
In this section, we formally define the scheduling problem, the algorithmic models under consideration, and the metrics used for performance evaluation.
@@ 2.1 Problem Formulation
A job becomes known to the algorithm at its release time $r_{j}$ and must be assigned to a machine at or after $r_{j}$. The job must be processed on a single machine at a time, and must be processed for $p_{j}$ time in total to complete. One machine can serve at most one job at a time.
We introduce standard notation to describe the status of a job or a machine at a given time $t$. A machine is said to be busy at time $t$ if it is processing a job, and idle otherwise. Time $t$ is said to be busy if all machines are busy at time $t$. A job $j$ is said to be waiting at time $t$ if it has been released but not yet started, i.e., $r_{j}\leq t<s_{j}$. A job is said to be processing at time $t$ if it is being processed by a machine.
@@ 2.2 Scheduling Models
A schedule $\S$ assigns a start time $s_{j}\geq r_{j}$ to each job $j$ on one of the $m$ machines. This paper focuses on two primary models:
@@
  • Kill-and-Restart: Algorithms are permitted to terminate (kill) a job that is currently running. The job then returns to the set of waiting jobs and can be scheduled again (restart) from the beginning at a later time, possibly on a different machine. All work done prior to the termination is lost.
@@
  • Preemptive (without Migration): The algorithm is allowed to interrupt a job and later resume it on the same machine without losing any processed work.
  • Preemptive with Migration: The algorithm is allowed to interrupt a job and later resume it on any machine without losing any processed work.
2.3 Objective and Performance Analysis
Let $\S=\textup{{ALG}}(J)$ be the schedule produced by some algorithm ALG on job set $J$. We write $s_{j}(\S)$ for the start time of job $j$ in schedule $\S$, and define its completion time as $C_{j}(\S)=s_{j}(\S)+p_{j}$. The flow time of job $j$ is $F_{j}(\S)=C_{j}(\S)-r_{j}$, and the objective is to minimize the total flow time on the job set $J$:
$F(\S)=\sum_{j\in J}F_{j}(\S).$
When the schedule $\S$ is clear from context, we omit it from the notation and write $s_{j}$, $C_{j}$, and $F_{j}$ instead.
@@ 2.3 Objective and Performance Analysis
The performance of an online algorithm is measured by its competitive ratio. An algorithm ALG is $\Gamma$-competitive [^4] The competitive ratio, $\Gamma$, is often a function $\Gamma(n,m)$ of the number of jobs $n$ and machines $m$ in this paper. if for any job instance $J$:
$F(\textup{{ALG}}(J))\leq\Gamma\cdot F(\textup{{OPT}}(J))$
where $\textup{{OPT}}(J)$ is the total flow time of an optimal offline solution. For the purpose of analysis, we define OPT differently for our upper and lower bounds:
  • For our upper bounds, we compare our online algorithm against a more powerful adversary: OPT is the optimal offline preemptive schedule without migration.
  • For our lower bounds, we use the offline non-preemptive schedule as OPT.
For convenience, when the context is clear, we will use OPT and ALG to denote the costs $F(\textup{{OPT}}(J))$ and $F(\textup{{ALG}}(J))$, respectively.
@@ 3.1 The Challenge of Dynamic Re-Classification
To achieve this, we maintain a set of active large jobs, denoted $A$, containing the $\ell$ jobs with the largest processing times seen so far. The logic of our framework defines a clear life cycle for every job based on its interaction with this set.
@@
      Remove $j$ from $A$ and add $i$ to $A$;
@@
       // The scheduler will actually process $j$ when it processes $j^{\prime}$.
@@ 3.2 Handling Small Jobs with Non-Preemptive Shortest Job First
Lemma 2 demonstrates the power of this partition by bounding the flow time contribution from this large group of small jobs. For example, when we set the rank threshold $\ell=\sqrt{n}$, our framework ensures that all small jobs have a processing time $\tau\leq\textup{{OPT}}/\sqrt{n}$. For clarity, we will first prove the lemma as stated. In Appendix A, we prove a more general version of this result (Lemma 29) that is robust to interference from large jobs, where machines may be subject to initial “blocking periods.”, during which they are unavailable for scheduling small jobs. Specifically, given a vector of initial blocking times $\vec{b}$ that NSJF cannot use a machine $i$ in its blocking period $0,b_{i})$, we show that NSJF remains competitive with respect to the optimal flow time on an instance without blocking, up to an extra additive factor depending on the total blocking length. This general version is a fundamental component in the final analysis of our main algorithms.
@@
Proof of Lemma 2.
To begin, we introduce notation to compare the progress of NSJF against OPT. For any processing time threshold $p$ and time $t$, let $J_{\leq p}(t)$ be the set of jobs with $p_{j}\leq p$ and release time $r_{j}\leq t$. We measure progress on this set differently for each algorithm:
  • For OPT, let $V_{\leq p}(t)$ denote the total volume of jobs in $J_{\leq p}(t)$ that has been processed by time $t$.
  • For NSJF, let $V^{\prime}_{\leq p}(t)$ denote the total volume of jobs in $J_{\leq p}(t)$ that have started by time $t$.
Our first step is to prove the following volume-matching lemma.
Lemma 3.
For every $t\geq 0$, and every upper bound $p$, we have: $V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t).$
@@
Fix an arbitrary $p$, let $t_{0}$ be the last time some machine is idle by NSJF before $t+\tau$. Define $s_{j}$ as the latest start time, within the interval $(t_{0},\,t+\tau)$, at which NSJF schedules a job with $p_{j}>p$. If no such job exists, we set $s_{j}=t_{0}$ and conceptually assume that an infinitely large job starts at time $s_{j}$.
@@
We now analyze two cases to conclude the lemma. If $s_{j}\geq t$. We straightforwardly have:
@@
Otherwise, if $s_{j}<t$, the time interval $s_{j}+\tau,t+\tau)$ must be fully used by NSJF to process jobs with size at most $p$ (if any). We have:
@@
Denote $N_{\leq p}(t)$ (or $N^{\prime}_{\leq p}(t)$ below) as the number of jobs in $J_{\leq p}(t)$ completed by time $t$ by OPT (or NSJF, respectively). Note that without the subscription $\leq p$, we remove the condition on job size. We prove:
Lemma 4.
For every $t\geq 0$, and every upper bound $p$, we have $N^{\prime}_{\leq p}(t+2\tau)\geq N_{\leq p}(t)$.
@@
We discuss two timepoints $t+\tau$ and $t+2\tau$. By Lemma 3, we have the volume bound: $V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t)$ for every $p$.
We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t+\tau$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$. That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all rank $i$.
To prove this, assume for contradiction that there exists the smallest rank $i$ such that $p^{\prime}_{i}>p_{i}$. Under this assumption, we have the following bound for OPT:
$V_{\leq p_{i}}(t)\geq\sum_{j=1}^{i}p_{j},$
since at least $i$ jobs of size at most $p_{i}$ are completed by time $t$ in OPT.
Now consider NSJF. Since the $i$-th smallest started job under NSJF has a size strictly greater than $p_{i}$, all started jobs of size $\leq p_{i}$ must be among the first i1i-1 jobs. Therefore, we have:
@@
where the second inequality uses the assumption that $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t+\tau)<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 3. Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all $i$.
Finally, this implies that the number of started jobs with $p_{j}\leq p$ by time $t+\tau$ under NSJF is at least the number of completed jobs with $p_{j}\leq p$ by time $t$ under OPT. Since all jobs started by time $t+\tau$ must be completed by time $t+2\tau$, it completes the proof of the lemma.
@@
Let $u$ denote the maximum completion time under NSJF. The flow times of NSJF and OPT can be expressed as:
$\textup{{OPT}}=\int_{0}^{\infty}\left(|J(t)|-N(t)\right)\,dt,\quad F(\textup{{NSJF}})=\int_{0}^{u}\left(|J(t)|-N^{\prime}(t)\right)\,dt,$
@@ 4.1 Single-Machine Case
   Classify job $i$ large or small based on Partition($i,\ell$) with $\ell=\lfloor\sqrt{n}\rfloor$ ;
@@
There exists an online polynomial time randomized non-preemptive algorithm for total flow time minimization on the single-machine case, e.g., Algorithm 2, is $O(\sqrt{n})$-competitive, against the preemptive offline solution.
@@
$F(\S_{1})\leq O(\sqrt{n})\cdot F(\textup{{OPT}}(S)).$
Moreover, since $S$ is weaker than $J$ by Lemma 1, we further have:
$F(\S_{1})\leq O(\sqrt{n})\cdot F(\textup{{OPT}}(S))\leq O(\sqrt{n})\cdot\textup{{OPT}}.$ (1)
@@
Let $\hat{F}_{j}$ denote the self-delay of $j$. We define:
@@
where $C^{\prime}_{j}$ is the completion time of $j$ in $\S_{2}$ immediately after its insertion (in the last update round of our algorithm). In Lemma 7, we show:
@@
For each large job $j$, the additional flow time arising from its delay of other previous jobs is bounded by $O(\sqrt{n})\cdot p_{j}$ in expectation, i.e., $\mathbb{E}(\delta_{j})\leq O(\sqrt{n})\cdot p_{j}$.
@@
Let $D_{k}$ be the set of jobs delayed by job $j$ when $w_{j}=k$. We aim to show that $\mathbb{E}[|D_{k}|]=O(\sqrt{n})$. The key claim is that for each job $i\neq j$, it appears in at most one $D_{k}$. This is because there is exactly $p_{j}$ idle time between any two adjacent insertion locations, since we collect $p_{j}$ units of idle time between them. This prevents a single job from being delayed by more than one possible insertion position of $j$.
Therefore, we have:
$\mathbb{E}[|D_{k}|]=\sum_{k=1}^{\sqrt{n}}|D_{k}|\cdot\Pr[w_{j}=k]\leq n\cdot\frac{1}{\lfloor\sqrt{n}\rfloor}\leq O(\sqrt{n}).$
@@
For a fixed large job $j$, the self-delay is either $r_{j^{\prime}}-r_{j}$, where $j^{\prime}$ is the proxied virtual job, or $C^{\prime}_{j}-r_{j}$ if $j$ is unproxied. Both types of self-delay arise from only two sources: collecting idle time or passing through busy time.
First, consider the idle time collected. Since $w_{j}\leq\sqrt{n}$, the total idle time accumulated for each large job $j$ is at most $\sqrt{n}\cdot p_{j}$. Summing over all large jobs, we have:
$\sum_{j\in L}\text{(idle contribution to self-delay)}\leq\sqrt{n}\cdot\sum_{j\in L}p_{j}\leq\sqrt{n}\cdot\textup{{OPT}}.$
Next, we analyze the busy time passed through by jobs. This may occur during the intervals $r_{j^{\prime}},r_{j})$ (for proxied jobs) or $r_{j},C^{\prime}_{j})$ (for unproxied jobs). In both cases, the large job remains unproxied throughout the interval. By Lemma 1, the number of unproxied large jobs at any time $t$ is at most $\sqrt{n}$. Since the total length of busy intervals is at most the total processing time of all jobs, we have:
$\sum_{j\in L}\text{(busy-time contribution to self-delay)}\leq\sqrt{n}\cdot\textup{{OPT}}.$
Combining both contributions, the total self-delay across all large jobs is bounded by

### v3

@@ 1.4 Paper Organization
The remainder of this paper is organized as follows. In Section 3, we introduce our core technical contribution: the online rank-based partitioning framework that serves as the foundation for all our algorithms, and the analysis of the NSJF algorithm for handling small jobs within this framework. In the deterministic algorithms, we use a generalized analysis of the NSJF algorithm, which is detailed in Appendix A.
We then present our main algorithmic results. Section 4 develops our $O(\sqrt{n/m})$-competitive randomized non-preemptive algorithm. Section 5 presents our deterministic non-preemptive $O(n/m^{2}+\sqrt{n/m}\log m)$-competitive algorithm. Section 6 details our deterministic $O(\sqrt{n/m})$-competitive algorithm using kill-and-restart for $m\geq 2$, and then we extend it to the unknown-$n$ setting in Section 7.
@@ 1.5 Related Works
In the classic non-preemptive offline setting, the problem is known to be NP-hard. A key result by [25] provided an $O(\sqrt{n/m}\log(n/m))$-approximation algorithm, which stood as the best-known guarantee for many years. On the approximation hardness front, [23] established a polynomial-time lower bound of $\Omega(\sqrt{n})$ for the single-machine case, later complemented by [25], who presented an $\Omega(n^{1/3-\varepsilon})$ lower bound for multiple machines when $n\geq m^{4/\varepsilon}$. In the online setting, the problem is notoriously difficult, with a simple greedy algorithm achieving a $\Theta(n/m)$ competitive ratio, which for a long time was the best-known result for deterministic algorithms.
@@
To circumvent the hardness of non-preemptive scheduling, a significant body of work has focused on the preemptive model, where jobs can be suspended and resumed. In this setting, the folklore SRPT (Shortest Remaining Processing Time) policy is optimal on a single machine. For multiple machines, SRPT achieves a tight competitive ratio of $\Theta(\log(n/m))$ when job migration is allowed [25]. For the more restrictive model without migration, [1] presented an algorithm with a competitive ratio of $O(\log(\min\{\log n,\log P\}))$.
@@
Another prominent line of research has explored models that grant additional power to the online non-preemptive algorithm. One common approach is resource augmentation, where the algorithm is provided with faster or more machines than the adversary. [28] showed that an $O(1)$-competitive ratio is achievable with $m\log P$ machines, where $P$ is the max-to-min job size ratio. Another popular relaxation is the rejection model, where the algorithm can choose to reject a small fraction of jobs. [7] first introduced this notion for flow time, with subsequent works [17, 26] achieving an $O(1/\varepsilon^{3})$-competitive algorithm for rejecting an $\varepsilon$-fraction of jobs.
@@
The kill-and-restart capability, introduced by [29] in the context of makespan minimization, represents a minimal form of preemption. This model has been successfully applied to other objectives, demonstrating its power to surpass non-preemptive lower bounds. For minimizing total completion time, [30] designed a $3/2$-competitive algorithm, breaking the non-preemptive barrier of $1.582$. Even more dramatically, for non-clairvoyant scheduling, [20] showed that kill-and-restart reduces the competitive ratio for total completion time from $\Omega(n)$ to $O(1)$. Despite these successes, the potential of kill-and-restart for flow time minimization has remained largely unexplored. The primary work in this specific area was by [9], who established several lower bounds, including an $\Omega(\sqrt{n})$ bound for randomized algorithms on a single machine, but no upper bounds were known.
@@ 2.1 Problem Formulation
A job becomes known to the algorithm at its release time $r_{j}$ and must be assigned to a machine at or after $r_{j}$. Each job must be processed on a single machine at a time and requires $p_{j}$ units of processing to complete. Each machine can process at most one job at a time.
We introduce standard terminology for describing the state of jobs and machines. A machine is said to be busy at time $t$ if it is processing a job, and idle otherwise. Time $t$ is said to be busy if all machines are busy at time $t$. A job $j$ is said to be waiting at time $t$ if it has been released but is not currently being processed by any machine.
@@ 2.2 Scheduling Models
This paper focuses on two primary models:
@@
  • Kill-and-Restart: The algorithm may terminate (kill) a job that is currently running. The job returns to the pool of available jobs and can be started again from the beginning at a later time, possibly on a different machine. All work done prior to the termination is lost.
@@
  • Preemptive (without Migration): The algorithm may interrupt a job and later resume it on the same machine without losing any processed work.
  • Preemptive with Migration: The algorithm may interrupt a job and later resume it on any machine without losing any processed work.
2.3 Objective and Performance Metric
We define $C_{j}$ as the time when job $j$ finishes its last unit of processing and its flow time as $F_{j}=C_{j}-r_{j}$. For the non-preemptive and kill-and-restart models, we write $s_{j}$ for the start time of job $j$’s (final successful) execution, and we have $C_{j}=s_{j}+p_{j}$. The objective is to minimize the total flow time:
$F(\S)=\sum_{j\in J}F_{j}.$
@@ 2.3 Objective and Performance Metric
The performance of an online algorithm is measured by its competitive ratio. An algorithm ALG is $\Gamma$-competitive [^4] The competitive ratio $\Gamma$ is often a function $\Gamma(n,m)$ of the number of jobs $n$ and machines $m$ in this paper. if for any job instance $J$:
$F(\textup{{ALG}}(J))\leq\Gamma\cdot F(\textup{{OPT}}(J)),$
where $\textup{{OPT}}(J)$ is the total flow time of an optimal offline solution. We define OPT differently for upper and lower bounds:
  • For upper bounds, OPT is the optimal offline preemptive schedule without migration.
  • For lower bounds, OPT is the optimal offline non-preemptive schedule.
For convenience, we use OPT and ALG to denote the costs $F(\textup{{OPT}}(J))$ and $F(\textup{{ALG}}(J))$, respectively, when the context is clear.
@@ 3.1 The Challenge of Dynamic Re-Classification
To achieve this, we maintain a set of active large jobs, denoted $A$, containing up to $\ell$ jobs with the largest processing times seen so far. The logic of our framework defines a clear life cycle for every job based on its interaction with this set.
@@
    // $j$ will be retired.
      Remove $j$ from $A$ and add $i$ to AA;
@@
       // The scheduler will actually process $j$ when processing $j^{\prime}$.
@@ 3.2 Handling Small Jobs with Non-Preemptive Shortest Job First
Lemma 2 demonstrates the power of this partition by bounding the flow time contribution from this large group of small jobs. For example, when we set the rank threshold $\ell=\sqrt{n}$, our framework ensures that all small jobs have a processing time $\tau\leq\textup{{OPT}}/\sqrt{n}$. For clarity, we will first prove the lemma as stated. In Appendix A, we prove a more general version of this result (Lemma 29) that is robust to interference from large jobs, where machines may be subject to initial “blocking periods,” during which they are unavailable for scheduling small jobs. Specifically, given a vector of initial blocking times $\vec{b}$ such that NSJF cannot use a machine $i$ in its blocking period $0,b_{i})$, we show that NSJF remains competitive with respect to the optimal flow time on an instance without blocking, up to an extra additive term depending on the total blocking length. This general version is a fundamental component in the final analysis of our main algorithms.
@@
For any processing time threshold $p$ and time $t$, let $J_{\leq p}(t)$ be the set of jobs with $p_{j}\leq p$ and release time $r_{j}\leq t$.
Step 1: Volume Matching.
For each algorithm, we track the volume of work associated with jobs in $J_{\leq p}(t)$. Let $V_{\leq p}(t)$ denote the total volume of work processed on jobs in $J_{\leq p}(t)$ by time $t$ under OPT, and let $V^{\prime}_{\leq p}(t)$ denote the total volume of jobs in $J_{\leq p}(t)$ that have been started by time $t$ under NSJF.
Lemma 3 (Volume Matching).
For every $t\geq 0$ and every threshold $p$, we have
$V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t).$
@@
Fix an arbitrary $p$, and let $t_{0}$ be the last time some machine is idle under NSJF before $t+\tau$. Let $j$ be the job with $p_{j}>p$ that is started latest by NSJF within the interval $(t_{0},t+\tau)$. If no such job exists, we let $s_{j}=t_{0}$ for a fictitious job $j$ of infinite size.
@@
We now analyze two cases. If $s_{j}\geq t$, we immediately have:
@@
Otherwise, if $s_{j}<t$, then since $s_{j}$ is the latest start of a job with $p_{j}>p$ after the last idle time $t_{0}$, all processing in $s_{j}+\tau,t+\tau)$ is on jobs of size at most $p$, and no machine is idle in this interval. Therefore:
@@
Step 2: Completion Matching.
Let $N_{\leq p}(t)$ and $N^{\prime}_{\leq p}(t)$ denote the number of jobs in $J_{\leq p}(t)$ completed by time $t$ under OPT and NSJF, respectively. When the subscript $\leq p$} is omitted, we count all completed jobs regardless of size.
Lemma 4 (Completion Matching).
For every $t\geq 0$ and every threshold $p$, we have
$N^{\prime}_{\leq p}(t+2\tau)\geq N_{\leq p}(t).$
@@
For each rank $i$, let $p_{i}$ be the size of the $i$-th smallest job completed under OPT by time $t$, and let $p^{\prime}_{i}$ be the size of the $i$-th smallest job started under NSJF by time $t+\tau$. Let $k$ and $k^{\prime}$ denote the total number of such jobs, respectively. We claim that $k^{\prime}\geq k$ and $p^{\prime}_{i}\leq p_{i}$ for all $i\leq k$.
Assume for contradiction that $i\leq k$ is the smallest rank such that either $i>k^{\prime}$ or $p^{\prime}_{i}>p_{i}$. For OPT, at least $i$ jobs of size at most $p_{i}$ are completed by time $t$, so:
$V_{\leq p_{i}}(t)\geq\sum_{j=1}^{i}p_{j}.$
For NSJF, all started jobs of size $\leq p_{i}$ are among the first i1i-1 jobs (either because only i1i-1 jobs have started, or because the $i$-th smallest is strictly larger than $p_{i}$). Therefore:
@@
where the second inequality uses $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. This gives $V^{\prime}_{\leq p_{i}}(t+\tau)<V_{\leq p_{i}}(t)$, contradicting Lemma 3.
Therefore $k^{\prime}\geq k$ and $p^{\prime}_{i}\leq p_{i}$ for all $i\leq k$, which implies that for every threshold $p$, the number of jobs of size $\leq p$ started by NSJF by time $t+\tau$ is at least the number completed by OPT by time $t$. Since all jobs started by time $t+\tau$ are completed by time $t+2\tau$, the lemma follows. ∎
Step 3: Bounding the Flow Time.
@@
Let $u$ denote the maximum completion time under NSJF. The total flow times can be expressed as:
@@ 4.1 Single-Machine Case
   Call Partition($i,\ell$) with $\ell=\lfloor\sqrt{n}\rfloor$ to classify job $i$ and update related jobs;
@@
Algorithm 2 is an online polynomial-time randomized non-preemptive algorithm for total flow time minimization on a single machine that is $O(\sqrt{n})$-competitive against the preemptive offline solution.
@@
$F(\S_{1})=O(\sqrt{n})\cdot F(\textup{{OPT}}(S)).$
Moreover, since $S$ is weaker than $J$ by Lemma 1, $F(\textup{{OPT}}(S))\leq\textup{{OPT}}$, and thus:
$F(\S_{1})=O(\sqrt{n})\cdot\textup{{OPT}}.$ (1)
@@
Let $\hat{F}_{j}$ denote the self-delay of $j$:
@@
where $\hat{C}_{j}$ is the completion time of $j$ in $\S_{2}$ immediately after its insertion (in the last update round of our algorithm). In Lemma 7, we show:
@@
For each large job $j$, the additional flow time arising from its delay of other previous jobs is bounded by $O(\sqrt{n})\cdot p_{j}$ in expectation, i.e., $\mathbb{E}(\delta_{j})=O(\sqrt{n})\cdot p_{j}$.
@@
Let $D_{k}$ be the set of jobs delayed by large job $j$ conditioned on the randomness $w_{j}=k$. To show that $\mathbb{E}[|D_{w_{j}}|]=O(\sqrt{n})$, the key claim is that every job $i\penalty 10000\ (i\neq j)$ appears in at most one of $D_{1},D_{2},\dots,D_{\lfloor\sqrt{n}\rfloor}$. This is because when inserting job $j$, there is exactly $p_{j}$ idle time between any two adjacent insertion locations. This prevents a single job from being delayed by more than one possible insertion position of $j$. Therefore, we have:
$\mathbb{E}[|D_{w_{j}}|]=\sum_{k=1}^{\lfloor\sqrt{n}\rfloor}|D_{k}|\cdot\Pr[w_{j}=k]\leq n\cdot\frac{1}{\lfloor\sqrt{n}\rfloor}=O(\sqrt{n}).$
@@
For a fixed large job $j$, the self-delay is either $r_{j^{\prime}}-r_{j}$ (if proxied by $j^{\prime}$) or $\hat{C}_{j}-r_{j}$ (if unproxied). We decompose the self-delay into three sources: collecting idle time, passing through busy time, and the job’s own processing time $p_{j}$.
First, consider the idle time collected. Since $w_{j}\leq\sqrt{n}$, the total idle time accumulated for each large job $j$ is at most $\sqrt{n}\cdot p_{j}$. Summing over all large jobs:
$\sum_{j\in L}\text{(idle contribution)}\leq\sqrt{n}\cdot\sum_{j\in L}p_{j}\leq\sqrt{n}\cdot\textup{{OPT}}.$
Next, we analyze the busy time passed through. For each large job $j$, this occurs during the interval $r_{j},r_{j^{\prime}})$ (if proxied) or $r_{j},\hat{s}_{j})$ (if unproxied, where $\hat{s}_{j}$ denotes the start time of $j$ immediately after its insertion). In both cases, the large job remains unproxied throughout the interval. By Lemma 1, the number of active unproxied large jobs at any time $t$ is at most $\sqrt{n}$, and there is at most one committed unproxied large job at any time $t$. Since the total length of busy intervals is at most OPT:
$\sum_{j\in L}\text{(busy-time contribution)}\leq(\sqrt{n}+1)\cdot\textup{{OPT}}.$
Finally, the processing time contribution is $\sum_{j\in L}p_{j}\leq\textup{{OPT}}$. Combining all three: